16.1 Evaluate [[cz + 3y – x²dS where S is the portion z = 2 – 3y + x2 that lies over the triangle in the xy-plane with vertices (0
Answer #1

Cha Given that SS (2+3y=92) ds -o S: Z= 2-3y +² 4 triangl in ny plance is (0,0), (2,6), (2,-4) in 0 putting Z=2-39+ 71² then